OpenStudy (anonymous):

Rectangle has a width thats 3 less than 2/3 the length, the perimeter is 114. Find L and W...thanks. I know the 2L + 2W = 114, but I'm still stuck! Thanks!

OpenStudy (anonymous):

ok so 2L + 2W = 114 you agreed :) with the information you have what could be the value ow W in terms of L ?

OpenStudy (anonymous):

*of

OpenStudy (anonymous):

L is = to 2/3 -3, is that right?

OpenStudy (anonymous):

they say W is 3 less then 2/3 the length so if you want W you could use a formula thats include both. W = 2/3L -3 ...?

OpenStudy (anonymous):

Would that be 2L + 2(2/3L-3) = 114 then?

OpenStudy (anonymous):

And if so, how do you solve that?

OpenStudy (anonymous):

you distribute the *2 : 2L + (2*2/3L - 3*2) ?

OpenStudy (anonymous):

1.3L - 6 ?

OpenStudy (anonymous):

I'm super confused about how to solve this equation

OpenStudy (anonymous):

its trigonometric substitution :) so yeah distrubuting the 2 would do 2L + 4/3L - 6 = 114 now that you only have one stranger (sorry im french i dont know what you call them lol) you can isolate him :)

OpenStudy (anonymous):

I Will plug that in and attempt to solve. Merci!

OpenStudy (anonymous):

haha ca fais plaisir